# SAVeD: Learning to Denoise Low-SNR Video for Improved Downstream Performance

### v1

@@ 3.4.1 Pseudo-PSNR for Downstream Detection
$PSNR_{D_{KL}}=\frac{1}{N}\sum_{b\in{N}}D_{KL}(d_{b}\|\tilde{d}_{b})$ (5)
A visualization of this metric can be seen in Supplementary materials Fig. 11.
@@ 4 Experiments
We demonstrate that SAVeD can improve performance in low-SNR videos across medical and ecological applications (Sec. 5). We evaluate our denoised images on downstream tasks for detection, tracking, counting, and classification.
@@ 4.1 Datasets
Caltech Fish Counting 2022 (CFC22) [37] is designed for detection, tracking, and counting fish in low-signal-to-noise sonar video. This dataset contains 1,567 sonar videos from seven different cameras on three rivers in Alaska and Washington. The videos are grayscale, their resolutions range from 288x624 to 1,086x2,125, their frame rates range from 6.7 to 13.3 fps, and each video is on average 336 frames (38s) in duration [37]. In total, there are 527,215 frames with 8,254 unique fish, totaling 516k bounding boxes and 16.7 hours of video [37]. The dataset includes significant domain shifts (e.g., background topology, occlusion, fish densities, fish sizes, camera noise), requiring models to generalize effectively across varying conditions.
Breast Lesion Ultrasound Video Dataset [44] (BUV) is designed for detection and classification (benign or malignant) of breast lesions. The dataset contains 188 videos, of which 113 are malignant and 75 are benign. These videos collectively have 25,272 images, each with 1 detection; the number of ultrasound images in each video range from 28 to 413. Each video has a complete scan of the abnormal tissue. The dataset has a random train–test split of 150–38 videos respectively[44].
The Point-of-care Ultrasound dataset (POCUS) [8, 7] is a collection of convex and linear probe lung ultrasound images and videos to classify/diagnose COVID-19 and pneumonia. It contains 247 videos and 59 images from both convex and linear probes. We exclusively use the video portion of this set. There are 70 ultrasound videos showing COVID cases, 45 showing possible COVID, 51 videos of bacterial pneumonia, 6 videos of viral pneumonia, and 75 of healthy lungs. Videos are sampled at 10Hz (10 frames per second). We group frames by video as in Born et al. [8, 7]. In total, we extract 9,184 frames. The average width x height of the frames is 499 x 463 pixels.
@@
4.2 Training Procedure
We train SAVeD using the reconstruction objective in Section 3. During training, we rescale CFC22 and POCUS images to 1024x512 and BUV and Fluo images to 1024x1024. For POCUS, we use T of 0.1 seconds (10Hz), as that is what the downstream process uses. For all other datasets, we use all frames. For each dataset, we train over all splits. We train for 20 epochs for CFC22, 120 epochs for POCUS, 40 epochs for BUV, and 1000 epochs for Fluo; we found these numbers of epochs sufficient for training to converge. These took 20 hours, 0.5 hours, 2 hours, and 2 hours, respectively, on 2 RTX 4090 GPUs; this is less time than other network-based denoising methods as seen in Tab. 4 in the Supp Mat. Additional details, including hyperparameter configurations, are in the Supp Mat Sec. B.2.
After training SAVeD, we generate denoised frames for all splits. In the case of CFC22, we combine the denoised image as two channels and the background-subtracted frame, $(I_{v})_{t}-\bar{I_{v}}$, as the last channel. For POCUS, BUV, and Fluo, we combine the the denoised image as two channels and the median-filtered image as the last channel.
@@ 4.2 Training Procedure
4.3 Evaluation procedure
Given that none of our videos have clean (noise-free) versions, we use the downstream performance tasks’ metrics as proxies for our denoised performance. We also use our metric from Sec. 3.4.
Denoising for Detection, Tracking, and Counting. For CFC22, which has detection, tracking, and counting as downstream tasks, we follow a simplified version of the detection pipeline from Kay et al. [37] – we train a YOLOv5 model for 5 epochs with the longest side of an image set to 896 and no augmentations. We remove duplicate predictions using non-maximal suppression. We use mAP50 [22] to evaluate detection performance frame-by-frame. We use a pretrained-frozen ByteTrack tracker and calculate MOTA [6], HOTA [46], and IDF1 [54] scores for evaluation. More details and hyperparameter settings are in the Supplemental Materials Sec. B.3 and B.4. For counting, we use trajectories from the tracks to create nMAE scores, defined in Kay et al. [37], for each domain. The tracking and counting pipelines do not require training.
For BUV, we follow the training procedure of Lin et al. [44]; we also follow their final fine-tuning step and evaluation to generate an AP50 metric. Note that we know that breast lesions are darker spots in ultrasounds. As a result, we invert our reconstruction error to take the minimum positive difference rather than the maximum:
$\text{inv}S_{t,T}=\min(0,I_{t}-I_{t-T})+I_{t}+\min(0,I_{t}+I_{t+t})$ (6)
Denoising for Classification. For POCUS, we perform 5-fold cross-validation as in Born et al. [8, 7], ensuring that frames from the same video are all in the same fold. We use the fine-tuning strategy and hyperparameters from Born et al. [8, 7]. We calculate each class’s precision, recall, and F1 scores and then average the folds’ metrics to determine overall metrics.
@@ 5 Results
We find that SAVeD is able to accurately denoise objects of interest in low signal-to-noise video. It improves a range of downstream tasks in a way that is computationally less resource-intensive and yields higher performance than other denoising methods.
@@
SAVeD produces clear contiguous objects, where other methods do not, shown in Fig. 3 & 4.
@@
Fish Denoising: CFC22. SAVeD increases the contrast between fish and background, see Fig. 4. As such, the distributions of pixel intensities at the same location when fish are present and when they are not are distinct. This is shown in Tab. 1, where SAVeD’s $PSNR_{D_{KL}}$ is significantly higher than that of other methods.
Avg. KL-Divergence ($\uparrow$   )
@@
Table 1: $PSNR_{D_{KL}}$ Quantitative denoised methods KL-divergence metric between P(Fish) vs. Q (Non-fish) as distributions of pixel intensities. For all ground truth bounding boxes, P and Q are composed as follows: P – we take the set of pixels in each box from frames with objects. Q – we extract the set of pixels from the same box location from a frame where there is no object at that location. Raw=raw noisy frame $I_{t}$, CFC22++[37] = 3-channel image (raw, background-subtracted, frame-to-frame difference), Denoised(framewise)= denoised with $[\Phi,\Omega,\Psi]$ trained with $I_{t}$ as the target (i.e. no motion augmentation), SAVeD=denoised with motion augmentation, as in Sec. 3.2. We calculate the KL-divergence metric, discussed in Sec. 3.4.1. $\uparrow$ indicates the metric is better the larger it is. Best values are bolded, worst values are in italics.
Fluorescent Cells Denoising: Fluo.  We found that our denoising method increases the cells’ brightness relative to the background, as seen in Fig. 3. As is standard[59], and because the data size is small, we only perform qualitative analysis on Fluo.
@@ 5.2 Detection Performance
Fish Detection: CFC22. The detection performance of SAVeD denoised frames is better than detection performance of other denoised frames for CFC22. This is shown in Fig. 4 & 5 and Tab. 2. SAVeD improves detection performance in areas where objects and signal are rare. Our denoised frames result in an improvement of 43.2% and 9.4% test accuracy compared to the raw and background-subtracted frames respectively, and a 5.1% boost in performance compared to a three-channel image (raw, background-subtracted, and frame-to-frame-absolute difference) described as baseline++ in Kay et al. [37], but hereon referred to as CFC22++.
@@
Blind-Spot/Mask Networks
@@
Figure 5: Denoising improves detections where signal is infrequent. (a) the ground truth fish patch locations (from bounding box labels) normalized over the dataset; most fish pass by in the top region, fish crossings below are infrequent, thus there is more training signal in the top part of the videos. (b/c) patchwise detection performance of CFC22++[37] and SAVeD, repsectively, on the CFC22 dataset. Heatmaps indicate mAP50 performance over all frames of the test set at pixel patches. The more red a patch is, the higher the mAP50 of that patch; the more blue the patch is, the lower the mAP50. (d) the difference, SAVeD - CFC22++, with solid ellipses at regions of heightened performance and dashed ellipses around areas of lowered performance. Denoising improves detections in areas where signal is infrequent. On the other hand, detection performance declines in areas where signal is abundant. Additional patch maps can be seen in Fig. 8 in the Supp Mat.
Compared to the background-subtracted frames, there is a 10.5% reduction in error in the validation set and a 20.3% reduction in error on the test set. SAVeD reduces error by 5.76% and 14.5% compared to the CFC22++ frames on the validation set and test set respectively.
Breast Lesion Detection: BUV. SAVeD clarifies the breast lesion imagery, as seen in Fig. 3. As a result, it is significantly more accurate than other denoising methods on breast lesion detection. This is shown in Tab. 2.
@@ 5.3 Tracking and Counting Performance
Fish Tracking and Counting: CFC22. Compared to classical and other DNN-based denoising methods, frames denoised by SAVeD achieve higher downstream performance for tracking and counting. The distinction between fish and background in the denoised frames is stronger, leading to fewer false negatives and more true positives. Results can be seen in Fig. 6 and Tab. 2.
Figure 6: Quantitative tracking improvements through denoising. CFC22++ consists of the three-channel (background-subtracted, absolute-difference, raw) frames. The “Perfect Denoiser” refers to frames that have black backgrounds and white masks at the bounding box locations. Denoising results in higher MOTA scores for val and test; SAVeD boosts IDF1 and HOTA scores in test moreso than in val.
@@ 5.4 Categorization Performance
Lung Health Categorization: POCUS. Our method yields the best 5-fold cross-validation image classification score compared to classical and network-based denoising methods on lung categorization, shown in Tab. 2. Fig. 7 shows the precision-recall for the denoising methods on the Covid class – SAVeD has the highest accuracy for Covid classification. Additional per-class performance analysis is in Sec. A.2 of the Supplementary materials.
@@ 5.5 Ablations
Autoencoder vs. no Autoencoder. Using a DAE improves downstream detection performance over using the reconstruction targets alone for all targets. This can be seen in Tab. 3 and Tab. 5(d) in the Supplementary material.
Architectures Our small denoising architecture has better performance on downstream tasks compared to larger architectures. Comparisons of SAVeD with a vanilla Autoencoder, a UNet[56], and a UNet[56] with 3D convolutional kernels are shown in Tab. 2. For additional details and architectures, see Tab. 5 in the Supp. mat. The vanilla Autoencoder’s architecture is also explicitly defined in Tab. 7.
@@
Table 3: Effect of Different Motion Enhancements with and without the Denoising Network on CFC22. All detectors that leverage the DAE have superior performance to those that use only the motion-enhanced target on the test set. The modified signal is used as the reconstruction target for the denoising autencoder when it is present, and is the input signal for the downstream task when the autoencoder is not used. All results are on CNNs with skip connections with resolution 1024 and bottleneck 512.
6 Discussion
Future work. We are interested in training end-to-end: combining the representations from the denoiser and the downstream tasks. We are also interested in experimenting with more than 3 frames as input to broaden the motion signature. Finally, we recognize the shared qualities of each of these datasets and also understand that self-supervised methods are data-hungry [43, 61]. As such, one could explore the performance benefit of training on all datasets collectively to learn general low-SNR video properties.
7 Conclusion
We present SAVeD, a self-supervised denoising method that does not require noise-free video which improves downstream performance in low-SNR videos. This is based on the confirmed intuition that while there is motion in the foreground and background, their motion signatures are distinct, and a simple model can separate them to improve the signal-to-noise ratio. Our proposed method captures objects’ motion while leveraging autoencoders’ denoising capabilities to improve downstream task performance efficiently. Our approach is general and applicable to a range of low-SNR video tasks and domains.
Supplementary Material
Appendix A Additional Experimental Results
A.1 Additional CFC22 Ablation Results
As in the main paper, we evaluate CFC22 on the detection val/test splits, and show results using mAP50 across the dataset splits. We look at the effect of bottleneck size in the hourglass network, traditional augmentations, input resolution size, and reconstruction targets on how the trained denoiser affects downstream detection performance.
Traditional Augmentations. We apply salt-and-pepper noise, gaussian-blur, motion-blur, brightness, and erasing from the kornia.Augmentations library. We found that no traditional augmentations, though, improve downstream detection performance. Results can be seen in Tab. 5(c).
Raw frame ($I_{t}$) predicts the input frame alone.
Positive Frame Difference with current frame (PFDwTN). We discuss this in section 3.2. We experimented with T=2 (PFDwT2) and T=1 (PFDwT1), ultimately selecting T=1.
Standard Deviation across all frames (sigma) is taken across all of the frames loaded in a window of continuous frames, $\sigma(I_{t-N}:I_{t+N})$ where 2N+1 is the size of the window. We experimented with N=1 and N=2.
Sum frames minus N*background ($\Sigma-N\bar{I}$) sums all of the frames in a window size N and takes the positive difference $N*\bar{I}$ where $\bar{I}$ is the mean frame of all frames in a video: $\max(0,(\sum_{t}^{T}I_{t})-N\bar{I})$. We experimented with window sizes N=3 and N=5.
Visualizations of all of these can be seen in Fig. 9
CFC22 POCUS BUV
N2V[40] 12 days 0.75 hours 1.5 hours
UDVD[59] 8 days* 12 hours 23 hours
SAVeD 20 hours 0.5 hours 2 hour
Table 4: SAVeD is time-efficient. Note that UDVD took 8 days* to train CFC22, but UDVD trained CFC22 only for one epoch. For all other datasets, UDVD trained for 10 epochs all on 2 NVIDIA RTX 4090 GPUs.
@@ A.1 Additional CFC22 Ablation Results
(c) Augmentations. Augmentations appear to degrade performace. All augmentation experiments are named as $augmentation_{probability}$.
@@
(d) Reconstruction targets. Reconstruction targets including both the original frame and the next or previous frames do better than reconstruction targets incorporating information from just one. Reconstruction targets with the current frame in have *. All results are on CNNs with resolution 1024 and bottleneck 512 (with no SKIP connection).

### v2

@@ 3.4.1 Foreground-to-Background Divergence (FBD) Metric for Unsupervised Denoising
$FBD_{D_{KL}}=\frac{1}{N}\sum_{b\in{N}}D_{KL}(d_{b}\|\tilde{d}_{b})$ (5)
A visualization of this metric can be seen in Supplementary materials Fig. 12. A larger score (divergence) indicates greater distinguishability of objects from background.
@@ 4 Experiments
We demonstrate that SAVeD can improve performance in low-SNR videos across medical and ecological applications (Sec. 5). We evaluate our processed images on downstream tasks for detection, tracking, counting, and classification.
@@ 4.1 Datasets
Caltech Fish Counting 2022 (CFC22) [24] is designed for detection, tracking, and counting fish in low-signal-to-noise sonar video. This dataset contains 1,567 sonar videos from seven different cameras on three rivers in Alaska and Washington. The videos are grayscale, their resolutions range from 288x624 to 1,086x2,125, their frame rates range from 6.7 to 13.3 fps, and each video is on average 336 frames (38s) in duration [24]. In total, there are 527,215 frames with 8,254 unique fish, totaling 516k bounding boxes and 16.7 hours of video [24]. The dataset includes significant domain shifts (e.g., background topology, occlusion, fish densities, fish sizes, camera noise), requiring models to generalize effectively across conditions.
Breast Lesion Ultrasound Video Dataset [32] (BUV) is designed for classifying (benign or malignant) and localizing breast lesions. The dataset contains 188 videos, of which 113 are malignant and 75 are benign. These videos collectively have 25,272 images, each with 1 detection; the number of ultrasound images in each video range from 28 to 413. Each video has a complete scan of the abnormal tissue. The dataset has a random train–test split of 150–38 videos respectively[32].
The Point-of-care Ultrasound dataset (POCUS) [6, 5] contains convex and linear probe lung ultrasound images and videos for classifying COVID-19 and pneumonia. It includes 247 videos and 59 images; we use only the videos. Of these, 70 show COVID cases, 45 possible COVID, 51 bacterial pneumonia, 6 viral pneumonia, and 75 healthy lungs. Videos are sampled at 10Hz, and frames are grouped by video as in Born et al. [6, 5]. In total, we extract 9,184 frames with an average size of 499×463 pixels.
@@
4.2 Training Procedure – Denoiser
We train SAVeD using the reconstruction objective in Eq. 3. During training, we rescale CFC22 and POCUS images to 1024x512 and BUV and Fluo images to 1024x1024. For POCUS, we use a sample rate of 10Hz, as that is what the downstream process uses. For all other datasets, we use all frames. Because SAVeD is self-supervised, when training the denoiser for each dataset, we train over all data. We train for 20 epochs for CFC22, 120 epochs for POCUS, 40 epochs for BUV, and 1000 epochs for Fluo; we found these numbers of epochs sufficient for training to converge. These took 20 hours, 0.5 hours, 2 hours, and 2 hours, respectively, on 2 RTX 4090 GPUs; this is less time than other network-based denoising methods as seen in Tab. 1. Additional details, including hyperparameter configurations, are in the Supplemental Materials Sec. B.2.
After training SAVeD, we generate denoised frames for all splits. In the case of CFC22, we combine the denoised image as two channels and the background-subtracted frame, $(I_{v})_{t}-\bar{I_{v}}$, as the last channel. For POCUS, BUV, and Fluo, we combine the denoised image as two channels and the median-filtered image as the last channel.
@@ 4.2 Training Procedure – Denoiser
4.3 Evaluation procedure – Downstream Tasks
Given that none of our videos have clean (noise-free) versions, we use the downstream performance tasks’ metrics as proxies for our denoised performance. We use $FBD$ from Sec. 3.4 when detection annotations are available.
Denoising for Detection, Tracking, and Counting. For CFC22, which has detection, tracking, and counting as downstream tasks, we follow a simplified version of the detection pipeline from Kay et al. [24] – we train a YOLOv5 model for 5 epochs with the longest side of an image set to 896 and no augmentations. We remove duplicate predictions using non-maximal suppression. We use mAP50 [12] to evaluate detection performance frame-by-frame. We use a pretrained-frozen ByteTrack tracker and calculate MOTA [4], HOTA [34], and IDF1 [40] scores for evaluation. More details and hyperparameter settings are in the Supplemental Materials Sec. B.3 and B.4. For counting, we use trajectories from the tracks to create nMAE scores, defined in Kay et al. [24], for each domain. The tracking and counting pipelines are training-free.
For BUV, we follow the training procedure of Lin et al. [32]; we also follow their final fine-tuning step and evaluation to generate an AP50 metric. Note that we know that breast lesions are darker (rather than brighter) spots in ultrasounds. As a result, we invert our reconstruction error to take the minimum difference rather than the maximum:
$\text{inv}S_{t,T}=\min(0,I_{t+T}-I_{t})+I_{t}+\min(0,I_{t-T}-I_{t})$ (6)
Denoising for Classification. For POCUS, we perform 5-fold cross-validation following Born et al. [6, 5], ensuring all frames from a video remain in the same fold. We adopt their fine-tuning strategy and hyperparameters. For evaluation, we compute per-class precision, recall, and F1, then average across folds to obtain overall metrics.
@@ 5 Results
SAVeD boost signal of objects of interest in low signal-to-noise video. It improves a range of downstream tasks in a way that is computationally less resource-intensive and yields higher performance than other denoising methods.
Train-Time (hours) CFC22 POCUS BUV
N2V [26] 144 0.75 1.5
UDVD [44] 96* 12 23
UMVD [1] 91* 34 42
LG-BPN [54] 101 6 4
SAVeD (Ours) 20 0.5 1
Table 1: SAVeD is resource-efficient during training. Note that UDVD and UMVD (U*VD) took 8 days to train CFC22, but U*VD trained CFC22 only for one epoch. For all other datasets, U*VD both trained for 10 epochs on 2 NVIDIA RTX 4090 GPUs.
@@
SAVeD produces clear contiguous objects, where other methods do not, shown in Fig. 4 & 5. SAVeD is also able train to convergence in 22% of the time of the next-quickest network-based method – more can be seen in Tab. 1.
@@
Fish Denoising: CFC22. SAVeD increases the contrast between fish and background (Fig. 5). As such, the distributions of pixel intensities at the same location when fish are present and when they are not are distinct. This is shown in Tab. 2, where SAVeD’s $FBD$ is significantly higher than that of other methods.
FBDFBD ($\uparrow$   )
@@
Table 2: FBD Eq. 5 between P(Fish) vs. Q (Non-fish). For all ground truth bounding boxes, P and Q are composed as follows: P – we take the set of pixels in each box from frames with objects. Q – we extract the set of pixels from the same box location from a frame where there is no object at that location. Raw=raw noisy frame $I_{t}$, CFC22++[24] = 3-channel image (raw, background-subtracted, frame-to-frame difference), SAVeD=denoised with motion augmentation, as in Sec. 3.2. We calculate the KL-divergence metric, discussed in Sec. 3.4.1. $\uparrow$ indicates the metric is better the larger it is. Best values are bolded, worst values are in italics.
Fluorescent Cells Denoising: Fluo. SAVeD increases the cells’ brightness relative to the background, as seen in Fig. 4. As is standard[44], and because the data size is small, we only perform qualitative analysis on Fluo.
@@ 5.2 Detection Performance
Fish Detection: CFC22. The detection performance of SAVeD-processed frames is better than detection performance of other processed frames for CFC22. This is shown in Fig. 5 & 6 and Tab. 3. SAVeD improves detection performance in areas where objects and signal are rare. SAVeD frames result in an improvement of 43.2% and 9.4% test accuracy compared to the raw and background-subtracted frames respectively, and a 5.1% boost in performance compared to a three-channel image (raw, background-subtracted, and frame-to-frame-absolute difference) described as baseline++ in Kay et al. [24], but hereon referred to as CFC22++.
@@
Deep-Learning-Based Image/Video Denoising/Restoration
@@
Figure 6: Denoising improves detections where signal is infrequent. (a) the ground truth fish patch locations (from bounding box labels) normalized over the dataset; most fish pass by in the top region, fish crossings below are infrequent, thus there is more training signal in the top part of the videos. (b&c) patchwise detection performance of CFC22++[24] and SAVeD, respectively, on the CFC22 dataset. Heatmaps indicate mAP50 performance over all frames of the test set at pixel patches. The more red a patch is, the higher the mAP50 of that patch; the more blue the patch is, the lower the mAP50. (d) the difference, SAVeD - CFC22++, with solid ellipses at regions of heightened performance and dashed ellipses around areas of lowered performance. Denoising improves detections in areas where signal is infrequent. On the other hand, detection performance declines in areas where signal is abundant. Additional patch maps can be seen in Fig. 9 in the Supp Mat.
Breast Lesion Detection: BUV. SAVeD clarifies the breast lesion imagery, as seen in Fig. 4 and results in a boost of over 11% in breast lesion detection mAP50 scores compared to the next best method (UMVD). More is in Tab. 3.
@@ 5.3 Tracking and Counting Performance
Fish Tracking and Counting: CFC22. Compared to classical and other DNN methods, SAVeD frames yield higher tracking and counting performance (Fig. 7, Tab. 3). Stronger fish–background separation reduces false negatives and increases true positives in detections, subsequently improving tracker and count accuracy.
Figure 7: Quantitative tracking improvements. CFC22++ consists of the three-channel (background-subtracted, absolute-difference, raw) frames. The “Perfect Denoiser” refers to frames that have black backgrounds and white masks at the bounding box locations. Denoising results in higher MOTA scores for val and test; SAVeD boosts IDF1 and HOTA scores in test moreso than in val.
@@ 5.4 Categorization Performance
Lung Health Categorization: POCUS. Images processed through SAVeD yield the best 5-fold cross-validation image classification score compared to classical and network-based denoising methods on lung categorization, shown in Tab. 3. Fig. 8 shows the precision-recall for the denoising methods on the Covid class – SAVeD has the highest accuracy for Covid classification. Additional per-class performance analysis is in Sec. A.2 of the Supp. mat.
@@ 5.5 Ablations
Autoencoder vs. no Autoencoder. Using a DAE improves downstream detection performance over using the reconstruction targets alone for all targets. More detail is in Tab. 4 and Tab. 5(d) in the Supplementary material.
Architectures Our small architecture has better performance on downstream tasks compared to larger architectures. Comparisons of SAVeD with a vanilla Autoencoder, a UNet[42], and a UNet[42] with 3D convolutional kernels are shown in Tab. 3. For additional details and architectures, see Tab. 5 in the Supp. mat. The vanilla Autoencoder’s architecture is also explicitly defined in Tab. 7.
6 Discussion and Conclusion
We recognize that combining noisy frames adds to the noise of the overall signal rather than removing it. Work [3] has shown that denoising methods capture clean data’s underlying structure. Denoising autoencoders purposefully corrupt training data by adding noise or masking some of the input values [51, 17]. We rely on the autoencoder to remove noise implicitly by focusing on the largest reconstruction areas to minimize loss. This assumes that the objects of interest are larger than the noise signature. Indeed, we found that detection performance occasionally dipped for very small (e.g. $\leq$ 10 pixel) objects, when this assumption is not held.
We observe a change in frame noise when there is an object of interest vs. when there is no object of interest: namely, when there is no object of interest, the whole frame is bright, vs. when an object is in the frame, the object pops and the background becomes dim: ideally, the background would stay dim regardless. An approach to handle this would be adding a component to the loss to focus on background consistency.
Despite these challenges, SAVeD-processed frames outperform other denoising methods on downstream tasks while requiring fewer training resources.
Future work. We are interested in training end-to-end: combining the representations from the denoiser and the downstream tasks. We are also interested in exploring how to broaden and emphasize the motion signature. Finally, we recognize the shared qualities of each of these datasets and also understand that self-supervised methods are data-hungry [29, 46]. As such, one could explore the performance benefit of training on all datasets collectively to learn general low-SNR video properties.
Conclusion. We present SAVeD, a self-supervised denoising method that improves downstream performance in low-SNR videos without requiring clean data. Our approach is general and applicable to a range of low-SNR video tasks and domains. It is based on the confirmed intuition that while there is motion in the foreground and background, motion signatures between foreground and background are distinct, and a simple model can separate them to improve the SNR. We also introduce a new metric, $FBD$, to capture this relationship. SAVeD enhances object motion while leveraging autoencoders’ denoising capabilities to boost downstream performance efficiently.
Acknowledgments. This work was supported by the Caltech Resnick Sustainability Institute Impact Grant “Continuous, accurate and cost-effective counting of migrating salmon for conservation and fishery management in the Pacific Northwest.” We also thank Neehar Kondapaneni, Angela Gao, Sara Beery, Justin Kay, and Laure Delisle for helpful feedback.
Supplementary Material
Appendix A Additional Experimental Results
A.1 Additional CFC22 Ablation Results
As in the main paper, we evaluate CFC22 on the detection val/test splits, and show results using mAP50 across the dataset splits. We look at the effect of bottleneck size in the hourglass network, traditional augmentations, input resolution size, and reconstruction targets on how the trained denoiser affects downstream detection performance. We also look at the effect of downstream task performance when using the reconstruction target alone ($S_{t,T}$) compared with using the learned reconstruction ($\hat{S}_{t,T}$).
Traditional Augmentations. We apply salt-and-pepper noise, gaussian-blur, motion-blur, brightness, and erasing from the kornia.Augmentations library. We apply these augmentations when training the denoiser. We do not apply these augmentations when training downstream tasks. We found that no traditional augmentations to train the denoiser, though, improve downstream performance. Results can be seen in Tab. 5(c).
Raw frame ($I_{t}$) predicts the input (identity) frame alone.
Positive Frame Difference with current frame (PFDwTN). We discuss this in more detail in the main paper, Sec. 3.2. We experimented with T=2 (PFDwT2) and T=1 (PFDwT1), ultimately selecting T=1.
Standard Deviation across all frames ($\sigma$) is taken across all of the frames loaded in a window of continuous frames, $\sigma(I_{t-N}:I_{t+N})$ where 2N+1 is the size of the window. We experimented with N=1 and N=2.
Sum frames minus N*background ($\Sigma-N\bar{I_{v}}$) sums all of the frames in a window size N and takes the positive difference $N*\bar{I_{v}}$ where $\bar{I_{v}}$ is the mean frame of all frames in a video: $\max(0,(\sum_{t}^{T}I_{t})-N\bar{I_{v}})$. We experimented with window sizes N=3 and N=5.
Visualizations of all of these can be seen in Fig. 10
@@ A.1 Additional CFC22 Ablation Results
Table 4: Effect of Different Motion Enhancements with and without SAVeD’s Autoencoder Network (AE) on CFC22. All detectors that leverage the AE have superior performance to those that use only the motion-enhanced target on the test set. The modified signal is used as the reconstruction target for the denoising autencoder when it is present, and is the input signal for the downstream task when the autoencoder is not used. All results are on CNNs with skip connections with resolution 1024 and bottleneck 512.
@@
No augmentations 83.5 70.6 77.6
(c) Augmentations. Augmentations appear to degrade performace. All augmentation experiments are named as $augmentation_{probability}$. All networks are CNNs with skip connections with resolution 1024 and bottleneck 512.
@@
(d) Reconstruction targets. Reconstruction targets including both the original frame and the next or previous frames do better than reconstruction targets incorporating information from just one. Reconstruction targets with the current frame in have *. All results are on CNNs with resolution 1024 and bottleneck 512 with no skip connection.